# SeRL: Self-Play Reinforcement Learning for Large Language Models with Limited Data

### v1

@@ Appendix A Additional Related Works
Self-Iteration Methods. Multi-round iterative training enables models to reach their performance upper bound progressively, and self-iteration has been widely explored. Pang et al. (2024) constructs preference pairs using rule-based rewards for DPO training. Chen et al. (2024) adopts adversarial learning based on human-labeled responses, making the method reliant on label quality. Dong et al. (2024) performs DPO-based iteration with a fixed reward model, while Wu et al. (2024c) formulates training as a zero-sum game using a learned preference model. Rosset et al. (2024) proposes DNO, which uses an expert model to estimate win rates by regressing internal rewards through batch-policy iteration. In contrast, we propose a majority-vote reward mechanism that requires no external supervision. Compared to prior offline approaches (Yuan et al., 2025a; Zelikman et al., 2022; Wu et al., 2024b), our method adopts an online RL framework. We also address data-scarce scenarios by generating instructions through self-instruction that adapt to the evolving capabilities of the model, and by using a simple yet effective majority-vote-based reward for accurate and efficient self-iteration.
@@ Appendix C Implementation Details
In this section, we provide the implementation details of our experiments. All experiments are run on a cluster with 8× NVIDIA RTX A6000 GPUs, a 96-core Intel Xeon Gold 5318Y CPU, and 512 GB RAM. The source code will be made publicly available [^1] https://anonymous.4open.science/r/SERL-B731.
C.1 Prompt Template for Instruction Generation
We adapt the instruction generation prompt from prior self-instruction work (Wang et al., 2023) to better suit the mathematical domain. The modified prompt template is shown in Tab. C.1.
[htbp] Table C.1: Few-shot instruction generation prompt Please act as a professional math teacher.
@@ C.1 Prompt Template for Instruction Generation
C.2 SR-DPO Rewarding Prompt
Since the prompt used in the original SR-DPO paper (Yuan et al., 2025a) is designed for reward estimation on general tasks, it performs poorly on mathematical problems. To ensure a fair comparison in our experiments, we modify the original prompt to better suit the evaluation of mathematical tasks, as shown in Tab. C.2.
[htbp] Table C.2: Reward estimation prompt in mathematical tasks Review the user’s math question and the corresponding response using the additive 5-point scoring system described below. Points are accumulated based on the satisfaction of each criterion:
@@ C.2 SR-DPO Rewarding Prompt
C.3 RL Training Hyperparamters
The experimental settings (Hu et al., 2024) for training LLaMA-3.2-3B-Instruct and Qwen-2.5-7B-Instruct using our SeRL method are shown in Tab. 5 and Tab. 6.
@@ C.3 RL Training Hyperparamters
SeRL $\text{n}_{\text{votes}}$ = 16
@@
SeRL $\text{n}_{\text{votes}}$ = 16
@@
Appendix D Additional Reward Analysis
D.1 Reward Hacking
[…]
@@ D.1 Reward Hacking
To more clearly present the impact of the difficulty filtering strategy on generated instructions, we analyze the difficulty and diversity of the first-round dataset generated by LLaMA-3.2-3B-Instruct before and after the difficulty filtering strategy. We denote the generated dataset from iteration $n$ as $\mathcal{D}_{\text{gen}}^{n}=\bigcup_{t\in\mathcal{T}^{n}}\mathcal{B}_{\text{gen}}^{t}$, where $\mathcal{T}^{n}$ represents the set of all generation steps in iteration $n$. Accordingly, the first-round generated dataset is denoted as $\mathcal{D}_{\text{gen}}^{1}$. We hypothesize that more difficult instructions require more complex CoT reasoning, which tends to result in longer responses. Qwen-3-32B (Yang et al., 2025), with its strong overall performance, can correctly answer most questions in the dataset, making it a reliable proxy for assessing instruction difficulty. Therefore, we use Qwen-3-32B to answer the instructions in $\mathcal{D}^{1}_{\text{gen}}$ and compare response length distributions as a proxy for difficulty. As shown in Fig. 4(a) and Fig. 4(b), applying the difficulty filtering leads to a reduction in both extremely short and excessively long responses, aligning with our design goal of removing overly easy and overly difficult instructions. To assess data diversity, we compute the ROUGE-L similarity between each generated instruction $\boldsymbol{x}_{i}\in\mathcal{D}^{1}_{\text{gen}}$ and the $\mathcal{D}_{\text{seed}}$. For each $\boldsymbol{x}_{i}$, we define:
@@
Appendix E Additional Generated Instructions Analysis
@@
E.1 Length and Quality
We compare the length and quality score of the datasets $\mathcal{D}^{1}_{\text{gen}}$, $\mathcal{D}^{2}_{\text{gen}}$, $\mathcal{D}^{3}_{\text{gen}}$, $\mathcal{D}_{\text{seed}}$, and the MATH training set. The length is measured by the number of tokens obtained using the Qwen-3-32B tokenizer, where a longer instruction typically indicates higher complexity. Invalid long instructions composed of repetitive sentences have already been removed by our online filtering strategy. Among the remaining instructions, a greater average length suggests higher semantic and structural complexity. The quality score is computed by prompting Qwen-3-32B to evaluate each instruction based on criteria such as completeness of the instruction description, clarity, and overall usefulness or relevance. The evaluation prompt is provided in Tab. E.1.
[htbp] Table E.1: Instruction quality evaluation prompt You are an expert evaluator of math problems. Please evaluate the overall quality of the following math problem.
@@ E.2 Difficulty
E.3 Diversity
Verb-Noun Structures. We analyze the verb–noun structures of instructions in $\mathcal{D}^{1}_{\text{gen}}$. Specifically, we employ the Berkeley Neural Parser (Kitaev and Klein, 2018; Kitaev et al., 2019) to extract the main verbs nearest to the root and their first direct noun objects from each instruction. As illustrated in Fig. 5(b), we present the top 20 most frequent verbs along with their top 4 directly associated noun objects. The results reveal a diverse range of verb–noun constructions, highlighting the syntactic variety and richness of the generated instructions.
@@ E.3 Diversity
UMAP Projection. We project the BERT features of the instructions in $\mathcal{D}^{1}_{\text{gen}}$ using UMAP  (McInnes et al., 2020), as shown in Fig. 6. The visualization shows that the generated instruction points are not tightly clustered but instead exhibit a wide distribution similar to that of the MATH training set. This phenomenon further demonstrates the diversity of the generated dataset.
Appendix F Additional Analysis of the Self-Iteration Upper Bound
@@
As shown in Fig. 7, we observe that after multiple iterations, the Maj@16 performance of the model does not exhibit a consistent upward trend. Since Maj@16 serves as an empirical upper bound for Pass@1 under our framework, the lack of improvement in Maj@16 suggests that the performance ceiling of the model remains unchanged. This observation is consistent with the conclusion of Yue et al. (2025), which states that reinforcement learning with verifiable rewards primarily converts Pass@K performance into Pass@1 without fundamentally improving the capacity of the model. Although our method does not lead to incremental improvements in Maj@K, it still achieves substantial gains in Pass@1 performance in data-scarce scenarios through unsupervised reinforcement learning. These results highlight the practical utility of our approach, particularly in specialized, data-scarce domains where labeled data is limited.
Appendix G Broader Impact
This work explores reinforcement learning for enhancing the reasoning capabilities of LLMs in data-scarce scenarios, which has the potential to significantly broaden the accessibility and applicability of advanced AI systems. By reducing reliance on large-scale, high-quality labeled datasets, our method could enable the development of performant LLMs in data-scarce domains such as clinical diagnostics (Ullah et al., 2024; McDuff et al., 2025; Wang and Liu, 2025) and aerospace engineering (Liu et al., 2025a; Connolly, 2025; Yadav, 2024), where expert-annotated data is often expensive or impractical to obtain. However, the ability to generate and reinforce synthetic data also introduces potential risks. Without careful oversight, unsupervised self-improvement loops may propagate or even amplify biases present in the initial seed data. Additionally, models trained in data-scarce settings may exhibit spurious generalization patterns, especially in high-stakes domains like clinical decision-making or aerospace control. To mitigate these risks, we incorporate a robust online filtering mechanism to enforce baseline quality and control difficulty during data generation. Nevertheless, domain-specific filtering and safety measures will be essential in future deployments, especially when adapting the framework to sensitive or high-impact application areas.

### v2

@@ Appendix A Additional Related Works
Self-Iteration Methods. Multi-round iterative training enables models to reach their performance upper bound progressively, and self-iteration has been widely explored. Pang et al. [50] constructs preference pairs using rule-based rewards for DPO training. Chen et al. [3] adopts adversarial learning based on human-labeled responses, making the method reliant on label quality. Dong et al. [8] performs DPO-based iteration with a fixed reward model, while Wu et al. [70] formulates training as a zero-sum game using a learned preference model. Rosset et al. [52] proposes DNO, which uses an expert model to estimate win rates by regressing internal rewards through batch-policy iteration. In contrast, we propose a majority-voting reward mechanism that requires no external supervision. Compared to prior offline approaches [81, 84, 69], our method adopts an online RL framework. We also address data-scarce scenarios by generating instructions through self-instruction that adapt to the evolving capabilities of the model, and by using a simple yet effective majority-voting reward for accurate and efficient self-iteration.
@@ Appendix C Implementation Details
All experiments are run on a cluster with 8× NVIDIA RTX A6000 GPUs, a 96-core Intel Xeon Gold 5318Y CPU, and 512 GB RAM.
C.1 Details of Reinforce++ Algorithm
Assuming a training batch contains $n_{\text{bs}}$ questions and each question is associated with $n_{\text{vote}}$ sampled responses, $G_{ij}(s_{t},y_{t})$ denotes the return at the $t$-th token of the $j$-th sampled response $\mathbf{y}_{i,j}$ for the $i$-th question $\mathbf{x}_{i}$. We compute the mean and standard deviation of the return across the batch as:
$\text{mean}_{t}=\frac{1}{n_{\text{bs}}\cdot n_{\text{vote}}}\sum_{i=1}^{n_{\text{bs}}}\sum_{j=1}^{n_{\text{vote}}}G_{ij}(s_{t},y_{t}),\ \text{std}_{t}=\sqrt{\frac{1}{n_{\text{bs}}\cdot n_{\text{vote}}}\sum_{i=1}^{n_{\text{bs}}}\sum_{j=1}^{n_{\text{vote}}}\left(G_{ij}(s_{t},y_{t})-\text{mean}_{t}\right)^{2}}.$ (7)
Then, the advantage is computed as:
$\hat{A}_{ij}(s_{t},y_{t})=\frac{G_{ij}(s_{t},y_{t})-\text{mean}_{t}}{\text{std}_{t}}.$ (8)
We would like to further clarify the computation of $G_{ij}(s_{t},y_{t})$ as follows:
$G_{ij}(s_{t},y_{t})=\sum_{k=t}^{\mid\mathbf{y}_{i,j}\mid}R_{ij}(s_{k},y_{k})-KL_{ij}(k),$ (9)
where $R_{ij}(s_{k},y_{k})$ denotes the token-level reward at the $k$-th token of the $j$-th response to the $i$-th question and $KL_{ij}(k)$ denotes the token-level KL divergence at the $k$-th token between the current model and the initial model for the $j$-th response to the $i$-th question. A negative sign is applied to serve as a KL penalty. Reinforce++ adopts the $k_{1}$ estimator.
In practice, the reward is assigned at the response level, consistent with other related works [62, 89, 54]. Therefore, only the last token receives a reward, and the above equation 9 can be simplified as:
$G_{ij}(s_{t},y_{t})=R_{ij}(s_{T},y_{T})-\sum_{k=t}^{T}KL_{ij}(k)=R(\mathbf{x}_{i},\mathbf{y}_{j})-\sum_{k=t}^{T}KL_{ij}(k).$ (10)
Here, $R(\mathbf{x}_{i},\mathbf{y}_{j})$ denotes the reward assigned to the $j$-th response of the $i$-th question.
C.2 Details on Chain of Thought
During the RL sampling stage, we prompt the model with “think step by step and give the final answer”, encouraging it to first generate a chain of thought (CoT) and then produce the final answer.
C.3 Details on Filtering Strategies
The proposed online instruction filter consists of four components:
  • (1) The similarity filter computes the ROUGE-L score between each newly generated instruction and all existing instructions in the dataset. If the score exceeds a predefined threshold (set to 0.7 following the prior work [63]), indicating high similarity to existing instructions, the new instruction is filtered out to encourage diversity.
  • (2) The keywords filter removes instructions containing specific keywords such as "image", "graph", "picture", "file", "map", "draw", "plot", or "write a program", as they refer to visual content or capabilities beyond the model’s scope. In addition, instructions starting with punctuation or non-English characters are also excluded.
  • (3) The length filter removes instructions that are either excessively long or short. Instructions exceeding 150 words often contain redundant content or even include solutions, while those with fewer than 3 words typically lack sufficient context for problem solving. This filtering step helps maintain the overall quality and clarity of the generated instructions.
  • (4) The difficulty filter evaluates the proportion of majority answers among the generated responses for a given instruction. Instructions with excessively high or low majority proportions are filtered out to ensure that the retained instructions maintains suitable difficulty for model learning. We set the thresholds as $\gamma_{\text{diff}}=0.2$ and $\gamma_{\text{easy}}=0.8$, as shown in Tab. 6.
C.4 Details on Instruction Generation
We adapt the instruction generation prompt from prior self-instruction work [63] to better suit the mathematical domain. The modified prompt template is shown in Box C.1. The $n_{\text{shot}}$ examples used in the few-shot context are randomly sampled from both the seed data and the generated instructions. Specifically, one-fourth ($n_{\text{shot}}/4$) are drawn from the seed data, while the remaining three-fourths ($3n_{\text{shot}}/4$) from previously generated instructions.
Box C.1: Few-shot instruction generation prompt Please act as a professional math teacher.
@@ C.4 Details on Instruction Generation
C.5 SR-DPO Rewarding Prompt
Since the prompt used in the original SR-DPO paper [81] is designed for reward estimation on general tasks, it performs poorly on mathematical problems. To ensure a fair comparison in our experiments, we modify the original prompt to better suit the evaluation of mathematical tasks, as shown in Box C.2. Since it is also a self-rewarding method, we use the language model, which is to be trained, as the reward model to estimate the reward.
Box C.2: Reward estimation prompt in mathematical tasks Review the user’s math question and the corresponding response using the additive 5-point scoring system described below. Points are accumulated based on the satisfaction of each criterion:
@@ C.5 SR-DPO Rewarding Prompt
C.6 RL Hyperparameters
The experimental settings [20] for training LLaMA-3.2-3B-Instruct and Qwen-2.5-7B-Instruct using our SeRL method are shown in Tab. 6 and Tab. 7.
@@ C.6 RL Hyperparameters
SeRL $\text{n}_{\text{vote}}$ = 16
@@
SeRL $\text{n}_{\text{vote}}$ = 16
@@
Appendix D Additional Experiments
D.1 SFT Baselines
We conducted this experiment and compared SFT using the sampled truly correct responses of the model. As shown in Tab. 8, we observe a consistent drop in performance. This may be due to the limited quality of self-sampled responses or the model’s existing proficiency on those samples, resulting in minimal gains from further fine-tuning. To validate our findings, we additionally conduct SFT using data distilled from Qwen2.5-Math-7B-Instruct, and as we hypothesized, fine-tuning with higher-quality data is indeed necessary to enhance the capabilities of the model.
Table 8: Pass@1 comparison of different SFT strategies on LLaMA-3.2-3B-Instruct.
Methods MATH-500 MATH-Hard ASDiv College Math TabMWP
Initial 47.6 22.5 84.6 35.2 46.4
SFT on Majority responses 44.6 20.2 83.1 30.5 48.2
SFT on Correct responses 44.6 20.7 83.9 33.0 50.6
SFT on Distilled responses 49.0 23.6 87.1 36.4 66.1
D.2 Results of More Iterations
We additionally report results for 4 to 6 rounds in Tab. 9, where we observe that the performance of the model gradually converges. As discussed in Section 6, we believe this convergence behavior is largely due to the limited capacity of the underlying foundation model. This observation aligns with findings from EMPO [87], which suggests that pre-trained language models already possess strong reasoning capabilities. In this context, RL post-training may primarily help activate latent reasoning patterns learned during pretraining, rather than introduce new ones.
It is worth emphasizing that our goal is not to achieve unlimited iterative improvement through self-play, which is currently an unrealistic expectation and an unsolved problem across the field. As stated in the introduction, our work focuses on combining self-instruction and self-rewarding methods to incentivize the reasoning abilities of LLMs in data-scarce settings, aiming to achieve performance comparable to training on full high-quality data with verifiable rewards.
Table 9: Pass@1 results over additional iterations on mathematical benchmarks.
Models Methods MATH- 500 MATH- Hard ASDiv College Math TabMWP
LLaMA-3.2-3B-Instruct SeRL (iter4) 52.7 24.1 88.8 38.0 71.1
SeRL (iter5) 52.1 23.9 88.9 37.9 71.5
SeRL (iter6) 52.0 24.2 89.1 37.9 70.9
Qwen-2.5-7B-Instruct SeRL (iter4) 75.4 50.6 94.4 55.2 79.7
SeRL (iter5) 76.0 50.1 94.6 55.3 80.2
SeRL (iter6) 76.1 49.8 94.6 55.3 80.4
D.3 Additional Results in the Medical Domain
To demonstrate that our method can be applied to other domains with limited data, we also conducted experiments in the medical domain. Specifically, we used 500 randomly selected instructions from the MedQA [48] training set as the seed data for our method and compared it with a supervised reinforcement learning baseline trained on the full 10.2k dataset (RL-GT). We evaluated both models on three medical benchmarks: MedQA, PubMedQA [24], and NephSAP [68]. All evaluations are conducted using greedy decoding with a maximum of 1,024 new tokens. As shown in Tab. 10, our SeRL achieves performance comparable to the RL-GT baseline, demonstrating its effectiveness.
Table 10: Pass@1 performance comparison between our method and other baselines on the medical benchmark.
Models Methods MedQA PubMedQA NephSAP
LLaMA-3.2-3B-Instruct Initial 53.4 54.4 24.2
RL-GT 57.2 55.0 26.1
SeRL (Ours) 56.9 55.3 27.5
Qwen-2.5-7B-Instruct Initial 58.2 32.1 29.5
RL-GT 59.4 33.1 31.2
SeRL (Ours) 59.1 34.8 30.9
Appendix E Additional Reward Analysis
E.1 Comparison with Alternative Rewarding Methods
Self-rewarding methods are fundamentally designed to approximate the true reward. Their effectiveness can be evaluated by measuring how closely the estimated rewards align with the ground-truth rewards, as greater agreement indicates more accurate reward estimation. To quantify this, we calculate multiple metrics including cosine similarity, mean squared error (MSE), and mean absolute error (MAE) between the estimated rewards and the ground-truth rewards, where higher similarity and lower error indicate greater accuracy in reward estimation. In our evaluation, we treat the rule-based reward as the ground-truth reward and measure the alignment of the majority-voting, model-based, entropy-based [87], and CAI reward with it. CAI reward is the reward model we train based on the work of Bai et al. [2]. As shown in Tab 11, our method achieves the highest accuracy among all compared approaches.
Table 11: Similarity between different self-rewards and the ground-truth rewards on MATH500 using LLaMA3.2-3B-Instruct.
Methods Cosine ($\uparrow$) MAE ($\downarrow$) MSE ($\downarrow$)
Majority-voting reward (Ours) 0.75 0.30 0.60
Model-based reward 0.17 0.89 1.2
Entropy-based reward 0.65 0.45 0.65
CAI reward 0.01 1.0 1.41
E.2 Reward Hacking
As evidenced by the learning curves presented in Section 4.3, we observe that the training process becomes prone to reward hacking without the dual-end clipped difficulty filtering strategy. To further investigate the nature of this reward hacking behavior, we conduct a case study, as shown in Box E.1. Notably, we find that although the model correctly derives the answer $\left(3,\frac{\pi}{2}\right)$ during the Chain-of-Thought (CoT) reasoning, it still outputs “The final answer is: $\boxed{0}$” at the end. We also provide a detailed explanation in Section 4.3 of how the dual-end clipped difficulty filtering strategy helps prevent reward hacking by design.
Box E.1: Case study of LLaMA-3.2-3B-Instruct outputs after reward hacking. Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$ ## Step 1: Recall the conversion formulas from rectangular to polar coordinates.
The conversion from rectangular coordinates $(x,y)$ to polar coordinates $(r,\theta)$ can be done using the formulas: $r=\sqrt{x^{2}+y^{2}}$ and $\theta=\tan^{-1}\left(\frac{y}{x}\right)$.
## Step 2: Substitute the given coordinates.
Given the point $(0,3)$, substitute $x=0$ and $y=3$: $r=\sqrt{0^{2}+3^{2}}=3$, and $\theta=\tan^{-1}\left(\frac{3}{0}\right)$.
## Step 3: Evaluate the angle.
Since $\tan^{-1}(3/0)$ is undefined, note that the point lies on the positive $y$-axis, which corresponds to $\theta=\frac{\pi}{2}$.
## Step 4: Write the final answer.
The polar coordinates are: $\boxed{(3,\frac{\pi}{2})}$.
The final answer is: $\boxed{0}$
@@ E.2 Reward Hacking
To more clearly present the impact of the difficulty filtering strategy on generated instructions, we analyze the difficulty and diversity of the first-round dataset generated by LLaMA-3.2-3B-Instruct before and after the difficulty filtering strategy. We denote the generated dataset from iteration $n$ as $\mathcal{D}_{\text{gen}}^{n}=\bigcup_{t\in\mathcal{T}^{n}}\mathcal{B}_{\text{gen}}^{t}$, where $\mathcal{T}^{n}$ represents the set of all generation steps in iteration $n$. Accordingly, the first-round generated dataset is denoted as $\mathcal{D}_{\text{gen}}^{1}$. We hypothesize that more difficult instructions require more complex CoT reasoning, which tends to result in longer responses. Qwen-3-32B [76], with its strong overall performance, can correctly answer most questions in the dataset, making it a reliable proxy for assessing instruction difficulty. Therefore, we use Qwen-3-32B to answer the instructions in $\mathcal{D}^{1}_{\text{gen}}$ and compare response length distributions as a proxy for difficulty. As shown in Fig. 4(a) and Fig. 4(b), applying the difficulty filtering leads to a reduction in both extremely short and excessively long responses, aligning with our design goal of removing overly easy and overly difficult instructions. To assess data diversity, we compute the ROUGE-L similarity between each generated instruction $\boldsymbol{x}_{i}\in\mathcal{D}^{1}_{\text{gen}}$ and the $\mathcal{D}_{\text{seed}}$. For each $\boldsymbol{x}_{i}$, we define:
@@
Appendix F Additional Generated Instructions Analysis
@@
F.1 Length and Quality
We compare the length and quality score of the datasets $\mathcal{D}^{1}_{\text{gen}}$, $\mathcal{D}^{2}_{\text{gen}}$, $\mathcal{D}^{3}_{\text{gen}}$, $\mathcal{D}_{\text{seed}}$, and the MATH training set. The length is measured by the number of tokens obtained using the Qwen-3-32B tokenizer, where a longer instruction typically indicates higher complexity. Invalid long instructions composed of repetitive sentences have already been removed by our online filtering strategy. Among the remaining instructions, a greater average length suggests higher semantic and structural complexity. The quality score is computed by prompting Qwen-3-32B to evaluate each instruction based on criteria such as completeness of the instruction description, clarity, and overall usefulness or relevance. The evaluation prompt is provided in Box F.1.
Box F.1: Instruction quality evaluation prompt You are an expert evaluator of math problems. Please evaluate the overall quality of the following math problem.
@@ F.2 Difficulty
F.3 Diversity
Verb-Noun Structures. We analyze the verb–noun structures of instructions in $\mathcal{D}^{1}_{\text{gen}}$. Specifically, we employ the Berkeley Neural Parser [29, 28] to extract the main verbs nearest to the root and their first direct noun objects from each instruction. Fig. 5(b) shows the top 20 verbs with their top 4 associated nouns, revealing diverse and syntactically rich verb–noun constructions in the generated instructions.
@@ F.3 Diversity
UMAP Projection. We project the BERT features of the instructions in $\mathcal{D}^{1}_{\text{gen}}$ using UMAP [45], as shown in Fig. 6. The visualization shows that the generated instruction points are not tightly clustered but instead exhibit a wide distribution similar to that of the MATH training set. This phenomenon further demonstrates the diversity of the generated dataset.
Appendix G Additional Analysis of the Self-Iteration Upper Bound
@@
As shown in Fig. 7, we observe that after multiple iterations, the Maj@16 performance of the model does not exhibit a consistent upward trend. Since Maj@16 serves as an empirical upper bound for Pass@1 under our framework, the lack of improvement in Maj@16 suggests that the performance ceiling of the model remains unchanged. This observation is consistent with the conclusion of Yue et al. [83], which states that reinforcement learning with verifiable rewards primarily converts Pass@K performance into Pass@1 without fundamentally improving the capacity of the model. Although our method does not lead to incremental improvements in Maj@K, it still achieves substantial gains in Pass@1 performance in data-scarce scenarios through unsupervised reinforcement learning. These results highlight the practical utility of our approach, particularly in specialized, data-scarce domains where labeled data is limited.
Appendix H Broader Impact
This work explores reinforcement learning for enhancing the reasoning capabilities of LLMs in data-scarce scenarios, which has the potential to significantly broaden the accessibility and applicability of advanced AI systems. By reducing reliance on large-scale, high-quality labeled datasets, our method could enable the development of performant LLMs in data-scarce domains such as clinical diagnostics [58, 44, 60] and aerospace engineering [36, 5, 75], where expert-annotated data is often expensive or impractical to obtain. However, the ability to generate and reinforce synthetic data also introduces potential risks. Without careful oversight, unsupervised self-improvement loops may propagate or even amplify biases present in the initial seed data. Additionally, models trained in data-scarce settings may exhibit spurious generalization patterns, especially in high-stakes domains like clinical decision-making or aerospace control. To mitigate these risks, we incorporate a robust online filtering mechanism to enforce baseline quality and control difficulty during data generation. Nevertheless, domain-specific filtering and safety measures will be essential in future deployments, especially when adapting the framework to sensitive or high-impact application areas.